OpenStudy (anonymous):

Round your answer to the correct number of significant digits and show work. 1. How many grams of NaI would be used to produce a 2.0 M solution with a volume of 10.00 L? 2. What is the molarity of a 2.50 x 102 cm3 solution containing 450.0 g of C6H12O6 ? 3. How many moles of Sr(NO3)2 would be used in the preparation of 20.50 L of a 30.5 M solution?

OpenStudy (anonymous):

20/10.00 149.89X20=2997.8 not sure

OpenStudy (anonymous):

2) 450/250 =1.8

OpenStudy (anonymous):

3)30.5=625.25/20.50 unsure what to do next on that one

OpenStudy (anonymous):

@MALTMAN0004 2.0/10.00 149.89X2.0=30.g

OpenStudy (anonymous):

okay i see it

OpenStudy (aaronq):

If this "2.0/10.00 149.89X2.0=30.g" is for "1. How many grams of NaI would be used to produce a 2.0 M solution with a volume of 10.00 L?" then it's wrong. first find the moles: Molarity= n of solute/L of solution 2 M = n/10 L -> n =20 moles of NaI then n=m/M, (i.e. moles=mass/Molar mass) -> m=n*M m= 20*149.89 g/mol so the first post was correct, but it was a fluke :P

OpenStudy (anonymous):

he just didnt see the decimal. thank you though